# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 10 Apr 2026 v3

@@ 7 Oracle Complexity
This section turns the Fisher-geometry mean-square bounds of Section 6 into an explicit oracle-complexity guarantee for achieving $\varepsilon$-stationarity. The steps are standard—smoothness descent, averaged-gradient bound, martingale concentration (cf. Nesterov, 2004, Chapter 2)—but we express the result in the natural Fisher dual norm and keep the statistical conditioning explicit.
@@ 7.3 High-probability oracle complexity
Theorem 7.3 (Oracle complexity for Fisher-dual stationarity).
@@
In particular, choosing $(\eta,T)$ so that the right-hand side is $\leq\varepsilon^{2}$ implies $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}\leq\varepsilon$ with probability at least $1-\delta$.
Equivalently, in terms of oracle calls $N:=Tb$ (sample-gradient evaluations), one sufficient scaling (up to constants and ignoring the transient term $\frac{L(\theta_{0})-L(\theta^{\star})}{\eta T\lambda_{\min}(F^{\star})}$) is
$N\;=\;\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})}{\varepsilon^{2}}\,\log\!\frac{1}{\delta}\right).$ (46)
@@
Step 1 (descent inequality). By $M$-smoothness and $\eta\leq 1/(4M)$, expand $L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L,\theta_{t+1}-\theta_{t}\rangle+\frac{M}{2}\|\theta_{t+1}-\theta_{t}\|^{2}$ with $\theta_{t+1}-\theta_{t}=-\eta(\nabla L(\theta_{t})+\xi_{t})$. After expanding and using $\eta\leq 1/(4M)$, so $1-M\eta/2\geq 1-1/8=7/8$ and $\eta M\leq 1/4$:
$L(\theta_{t+1})\leq L(\theta_{t})-\tfrac{3\eta}{4}\|\nabla L(\theta_{t})\|_{2}^{2}+\eta(1+M\eta)\langle\nabla L(\theta_{t}),\xi_{t}\rangle+\tfrac{M\eta^{2}}{2}\|\xi_{t}\|_{2}^{2}.$
@@
Step 2 (define martingale). Let $\Delta M_{t+1}:=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ and $M_{T}:=\sum_{t=0}^{T-1}\Delta M_{t+1}$; since $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$, this is a martingale.
Step 3 (sum the descent). Summing from $t=0$ to T1T-1:
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}}{2}\,Q_{T},\quad\Delta L:=L(\theta_{0})-L(\theta^{\star}),$
where $Q_{T}:=\sum_{t<T}\|\xi_{t}\|_{2}^{2}$. To replace the random $Q_{T}$ by a deterministic bound we need concentration, not just conditional moments. Under conditional sub-Gaussianity, $\|\xi_{t}\|_{2}^{2}$ is conditionally sub-exponential with $\mathbb{E}[\|\xi_{t}\|_{2}^{2}\mid\mathcal{F}_{t}]\leq b^{-1}\mathrm{Tr}(F^{\star})$; Bernstein’s inequality for sums of conditionally sub-exponential random variables (see, e.g., Wainwright 2019, Cor. 2.21) gives, with probability at least $1-\delta$, $Q_{T}\leq(T/b)\mathrm{Tr}(F^{\star})+c\sqrt{T\,d}\,(\|F^{\star}\|_{\operatorname{op}}/b)\sqrt{\log(1/\delta)}.$ Since $M\eta^{2}/2\leq\eta/8$ (as $\eta\leq 1/(4M)$), the deviation is $O(\eta\sqrt{T/b}\,\sqrt{d}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)}\,/b)$, which is lower-order relative to the bias term $\Delta L/(\eta T)$ in the final bound. Absorbing it into the constant $C$ (and taking a union bound over the two $\delta$-events, i.e. replacing $\delta$ by $\delta/2$ throughout):
@@
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})\log\!\tfrac{2}{\delta},$
with probability at least $1-\delta$.
@@
Step 4 (sub-Gaussian martingale concentration). Each increment satisfies $\Delta M_{t+1}=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ where $\nabla L(\theta_{t})$ is $\mathcal{F}_{t}$-measurable. By the conditional sub-Gaussianity assumption, the scalar $\langle v,\xi_{t}\rangle$ with $v:=\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|_{2}$ has sub-Gaussian parameter $\sigma_{t}^{2}\leq\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $c_{t}^{2}:=\eta^{2}\|\nabla L(\theta_{t})\|_{2}^{2}\|F^{\star}\|_{\operatorname{op}}/b$. By the standard exponential supermartingale bound for sums of conditionally sub-Gaussian increments (see, e.g., Wainwright 2019, Theorem 2.19), with probability at least $1-\delta$, $M_{T}\leq\sqrt{2\sum_{t<T}c_{t}^{2}\,\log(1/\delta)}=\eta\sqrt{2(\|F^{\star}\|_{\operatorname{op}}/b)\,S_{T}\,\log(1/\delta)}$, where $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$.
Step 5 (solve for gradient bound). Substituting the Step 4 tail bound into Step 3 gives a quadratic inequality in $\sqrt{S_{T}}$; solving:
@@
$\min_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}\;\leq\;\frac{S_{T}}{T}\;\leq\;C\!\left(\frac{\Delta L}{\eta T}+\frac{\eta}{b}\mathrm{Tr}(F^{\star})\right)\!\log\!\frac{1}{\delta}.$
@@
establishing (45). The oracle complexity (46) follows by balancing the two terms and setting $N=Tb$. ∎
@@
Fix a total sampling budget $N$ and run $T=N/b$ iterations of constant-stepsize SGD with batch size $b$. Equating the bias term $\sim(L(\theta_{0})-L^{\star})/(\eta T\lambda_{\min}(F^{\star}))$ and the variance term $\sim\eta\kappa_{F}d_{\operatorname{eff}}/b$ in (45) yields the balanced stepsize $\eta^{\star}\propto\sqrt{b/(T\kappa_{F}d_{\operatorname{eff}})}$. Substituting $T=N/b$, the resulting Fisher-dual stationarity gap satisfies
@@
which is independent of $b$ at leading order. The practical implication for OR settings is that, under this budget-constrained regime, small batches lose nothing in worst-case guarantee while gaining more frequent iterates; the Fisher/Godambe geometry determines the constant but not the scaling in $N$.
@@ 8 Numerical Validation
We validate the theory through six diagnostic experiments summarized in Table 3. The experiments probe downstream consequences of the identified covariance $\mathrm{Cov}(g_{B})\approx b^{-1}G^{\star}$ (Theorem 4.3): if the identification is correct, the Lyapunov equilibrium, convergence rates, and directional variance structure must all agree with the predictions derived from $G^{\star}$. Appendix B additionally validates the covariance identity itself under misspecification and approximate exchangeability.
Table 3: Reproducibility capsule: key experimental parameters.
Experiment $d$ Replicates $\eta$ Steps Sweep variable
1 (Batch scaling) 2 500 0.05 200 000 $b\in\{1,2,4,\ldots,512\}$
2 (Angle sweep) 2 500 0.05 200 000 $\theta\in[0,\pi/2]$
3 (Rotation diagnostic) 2 $\phi\in[0,\pi/2]$ (theory)
4 (OU plateau, d=10d\!=\!10) 10 2 000 $0.5/\lambda_{\max}$ 300–10 000 $T$ (horizon)
5 ($1/N$ rate) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 $N=Tb$
6 (Directional var.) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 eigenvectors of $G^{\star}$
Experiments 1–3: Euler–Maruyama, $d=2$. Experiments 4–6: direct SGD recursion, $d=10$, $b=64$.
8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
The first claim is that the Fisher–Lyapunov balance (Corollary 5.7) is not merely a qualitative insight but a quantitatively accurate prediction of steady-state risk, and that the shape of the noise covariance—not just its scale—governs which directions receive noise energy at equilibrium. Three experiments validate this at increasing dimensionality.
Throughout, when the simulator uses finite-step (Euler–Maruyama or direct SGD) dynamics, the mathematically correct benchmark is the discrete Lyapunov equation (Proposition 5.9); the continuous-time OU prediction is its small-$\eta$ asymptotic reduction. The regime of approximation accuracy is governed by the product $\eta\lambda_{\max}(H^{\star})$: when this quantity is small (say $\leq 0.1$), the continuous and discrete benchmarks are nearly indistinguishable; for moderate stepsizes the discrete equation is the appropriate comparator.
8.1.1 Experiment 1: Fisher–Lyapunov Equilibrium and $1/b$ Scaling (OU Benchmark).
We begin with the local OU surrogate implied by the diffusion limit near a nondegenerate optimum:

### 15 Apr 2026 v4

@@ 7 Oracle Complexity
This section records a standard high-probability conversion of the Fisher-geometry mean-square bounds of Section 6 into an oracle-complexity guarantee for $\varepsilon$-stationarity in the Fisher dual norm. The proof is a textbook descent-lemma plus martingale concentration argument (cf. Nesterov, 2004; Wainwright, 2019, Ch. 2); the novelty is not the complexity theorem itself but the fact that, once the mini-batch covariance is identified as $b^{-1}F^{\star}$ (Theorem 4.3), the resulting constants read as intrinsic statistical quantities ($\kappa_{F}$, $d_{\mathrm{eff}}$) rather than as Euclidean conditioning.
@@ 7.3 High-probability oracle complexity
Theorem 7.3 (High-probability oracle complexity under the identified noise covariance).
@@
In particular, balancing the two terms on the right-hand side over $\eta$ yields
$\min_{0\leq t\leq T-1}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C^{\prime}\sqrt{\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\,(L(\theta_{0})-L(\theta^{\star}))}{N\,\lambda_{\min}(F^{\star})}}\,\log\!\frac{1}{\delta},$
where $N=Tb$. Hence the squared Fisher-dual stationarity gap decays as $N^{-1/2}$, and to achieve $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}\leq\varepsilon$ with probability at least $1-\delta$ it suffices that
$N\;=\;\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\,(L(\theta_{0})-L(\theta^{\star}))}{\lambda_{\min}(F^{\star})\,\varepsilon^{4}}\,\log^{2}\!\frac{1}{\delta}\right).$ (49)
The $\varepsilon^{-4}$ rate is the standard nonconvex-SGD rate for squared-gradient stationarity at level $\varepsilon^{2}$; it is not the $\varepsilon^{-2}$ rate of (strongly) convex stochastic optimization.
@@
Step 1 (descent inequality, correct sign). By $M$-smoothness, $L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\tfrac{M}{2}\|\theta_{t+1}-\theta_{t}\|^{2}$ with $\theta_{t+1}-\theta_{t}=-\eta(\nabla L(\theta_{t})+\xi_{t})$. Expanding $\|\nabla L(\theta_{t})+\xi_{t}\|^{2}=\|\nabla L(\theta_{t})\|^{2}+2\langle\nabla L(\theta_{t}),\xi_{t}\rangle+\|\xi_{t}\|^{2}$,
$L(\theta_{t+1})\;\leq\;L(\theta_{t})\;-\;\eta\bigl(1-\tfrac{M\eta}{2}\bigr)\|\nabla L(\theta_{t})\|^{2}\;-\;\eta(1-M\eta)\,\langle\nabla L(\theta_{t}),\xi_{t}\rangle\;+\;\tfrac{M\eta^{2}}{2}\|\xi_{t}\|^{2}.$
Under $\eta\leq 1/(4M)$, $1-M\eta/2\geq 7/8$ and $1-M\eta\geq 3/4$. Denote $\alpha:=1-M\eta\in[3/4,1)$.
@@
Step 2 (define martingale). Let $\Delta M_{t+1}:=-\eta\alpha\,\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ and $M_{T}:=\sum_{t=0}^{T-1}\Delta M_{t+1}$; since $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$, $(M_{t})$ is a martingale adapted to $(\mathcal{F}_{t})$. Write $\Delta L:=L(\theta_{0})-L(\theta^{\star})$, $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|^{2}$, and $Q_{T}:=\sum_{t<T}\|\xi_{t}\|^{2}$. Summing the Step-1 inequality from $t=0$ to T1T-1,
$\tfrac{7\eta}{8}\,S_{T}\;\leq\;\Delta L\;+\;M_{T}\;+\;\tfrac{M\eta^{2}}{2}\,Q_{T}.$ (50)
Step 3 ($Q_{T}$ tail via a Hanson–Wright-type bound). Under conditional sub-Gaussianity of $\xi_{t}$ with variance proxy $\Sigma_{t}\preceq b^{-1}F^{\star}$, the Hanson–Wright inequality for sub-Gaussian vectors (Vershynin, 2018, Thm. 6.2.1) gives, for an absolute constant $c_{0}$, $\Pr(\,|\|\xi_{t}\|^{2}-\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t}]|>s\mid\mathcal{F}_{t})\leq 2\exp(-c_{0}\min\{s^{2}b^{2}/\|F^{\star}\|_{F}^{2},\,sb/\|F^{\star}\|_{\operatorname{op}}\})$; so $\|\xi_{t}\|^{2}-\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t}]$ is conditionally sub-exponential with parameters $(\|F^{\star}\|_{F}^{2}/b^{2},\|F^{\star}\|_{\operatorname{op}}/b)$. Applying Bernstein’s inequality (Wainwright, 2019, Ch. 2) over $t<T$ gives, with probability at least $1-\delta/2$,
@@
$Q_{T}\;\leq\;\frac{T}{b}\,\mathrm{Tr}(F^{\star})\;+\;c_{1}\Bigl(\tfrac{\|F^{\star}\|_{F}}{b}\sqrt{T\log(2/\delta)}+\tfrac{\|F^{\star}\|_{\operatorname{op}}}{b}\log(2/\delta)\Bigr).$ (51)
Since $\|F^{\star}\|_{F}^{2}\leq d\,\|F^{\star}\|_{\operatorname{op}}^{2}$, the deviation term is $O(b^{-1}\sqrt{d\,T}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)})$, lower order than the leading $(T/b)\mathrm{Tr}(F^{\star})$ when $T\gtrsim d\log(1/\delta)$.
@@
Step 4 (two-sided martingale concentration). The scalar $\langle\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|,\,\xi_{t}\rangle$ is conditionally sub-Gaussian with parameter at most $\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $v_{t}^{2}:=\eta^{2}\alpha^{2}\,\|\nabla L(\theta_{t})\|^{2}\,\|F^{\star}\|_{\operatorname{op}}/b$. The standard two-sided Azuma–Hoeffding/Freedman-type bound for conditionally sub-Gaussian martingale sums (Wainwright, 2019, Thm. 2.19) gives, with probability at least $1-\delta/2$,
$|M_{T}|\;\leq\;\sqrt{2\sum_{t<T}v_{t}^{2}\,\log(4/\delta)}\;=\;\eta\alpha\,\sqrt{\frac{2\|F^{\star}\|_{\operatorname{op}}}{b}\,S_{T}\,\log(4/\delta)}.$ (52)
Step 5 (combine). On the intersection of the events in (51)–(52) (probability $\geq 1-\delta$ by union bound), substitute into (50) and use $|M_{T}|$ in place of $M_{T}$:
$\tfrac{7\eta}{8}\,S_{T}$ $\;\leq\;\Delta L\;+\;\eta\sqrt{\tfrac{2\|F^{\star}\|_{\operatorname{op}}}{b}\,S_{T}\,\log(4/\delta)}\;+\;\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})$
$\qquad\;+\;c_{2}\,\tfrac{M\eta^{2}\|F^{\star}\|_{\operatorname{op}}}{b}\Bigl(\sqrt{T\log(2/\delta)}+\log(2/\delta)\Bigr).$ (53)
The last term scales as $\eta^{2}\sqrt{T}/b$ (up to log factors), strictly lower order than the leading variance term $\eta^{2}T\mathrm{Tr}(F^{\star})/b$ once $T\gtrsim\log(1/\delta)$, and is absorbed into the constant $C$ below. Solving the resulting quadratic inequality in $\sqrt{S_{T}}$ for $S_{T}/T\geq\min_{t<T}\|\nabla L(\theta_{t})\|^{2}$ yields
@@
$\min_{t<T}\|\nabla L(\theta_{t})\|^{2}\;\leq\;C\!\left(\frac{\Delta L}{\eta T}+\frac{\eta}{b}\mathrm{Tr}(F^{\star})\right)\log(4/\delta),$
for a constant $C$ depending only on $M$, $c_{0}$, and $c_{1}$.
@@
establishing (48). For (49), balance the two terms by taking $\eta^{\star}=\sqrt{\Delta L\,b/\bigl(T\,\lambda_{\min}(F^{\star})\,\kappa_{F}\,d_{\operatorname{eff}}\bigr)}$, which gives the balanced RHS $\sim\sqrt{\Delta L\,\kappa_{F}\,d_{\operatorname{eff}}/(Tb\,\lambda_{\min})}\,\log(1/\delta)$, i.e., the squared-norm gap decays as $N^{-1/2}$ in $N=Tb$. Requiring this squared gap to be at most $\varepsilon^{2}$ yields the $\varepsilon^{-4}$ scaling of (49). ∎
@@
Fix a total sampling budget $N$ and run $T=N/b$ iterations of constant-stepsize SGD with batch size $b$. Balancing the bias term $\sim(L(\theta_{0})-L^{\star})/(\eta T\lambda_{\min}(F^{\star}))$ and the variance term $\sim\eta\kappa_{F}d_{\operatorname{eff}}/b$ in (48) yields the balanced stepsize $\eta^{\star}\propto\sqrt{b/(T\kappa_{F}d_{\operatorname{eff}})}$. Feasibility. The balanced stepsize must respect the theorem’s own step-size cap $\eta\leq 1/(4M)$. Substituting $T=N/b$ gives $\eta^{\star}=b\sqrt{\Delta L/(N\,\lambda_{\min}(F^{\star})\,\kappa_{F}\,d_{\operatorname{eff}})}$, so the cap $\eta^{\star}\leq 1/(4M)$ imposes the feasibility condition
$b\;\leq\;\sqrt{\frac{N\,\lambda_{\min}(F^{\star})\,\kappa_{F}\,d_{\operatorname{eff}}}{16\,M^{2}\,\Delta L}}.$
The batch-invariance conclusion below is a statement within this feasibility region (so $b$ may grow like $\sqrt{N}$, not linearly in $N$), not at arbitrary $b$. Substituting $T=N/b$ and taking $\eta=\eta^{\star}$ within the feasibility region, the resulting squared Fisher-dual stationarity gap satisfies
@@
so the norm gap decays as $N^{-1/4}$. Within the feasibility region the bound is independent of $b$ at leading order, which suggests—under a fixed oracle budget—that small batches do not degrade the worst-case guarantee while gaining more frequent iterates; the identified geometry determines the constant but not the scaling in $N$.
@@ 8 Numerical Validation
The empirical program is organized in three layers, each answering a distinct question about the identified covariance and its downstream consequences:
  1. (i)
    Direct identification (Experiment 1). Does $\mathrm{Cov}(g_{B}(\theta^{\star}))\approx b^{-1}G^{\star}(\theta^{\star})$ hold in an actual mini-batching setup—including under model misspecification, where curvature and noise geometry diverge?
  2. (ii)
    Fluctuation-scale geometric illustrations (Experiment 2). What does the Lyapunov structure implied by the identified covariance look like in a low-dimensional setting where individual eigendirections can be visualized?
  3. (iii)
    Raw-scale quantitative validations (Experiments 3–6). Do the original-iterate-scale predictions—discrete Lyapunov plateau at constant step, $1/N$ rate under Robbins–Monro schedules, directional amplification, and the fluctuation-scale $(\eta/b)$-collapse (signature of Theorem 5.7)—hold quantitatively in direct SGD on higher-dimensional problems?
Layer (i) is the direct check of Theorem 4.3. Layer (ii) is interpretive: it visualizes the fluctuation-scale Lyapunov geometry on the normalized time-changed OU process of Theorem 5.7, and is not a direct Monte Carlo verification of raw SGD. Layer (iii) carries the main empirical weight: Experiments 3–6 are direct SGD recursions on the original iterate scale, and are the quantitative counterparts of the Lyapunov, rate, directional, and scaling-collapse predictions derived in Sections 57. Additional robustness diagnostics (approximate exchangeability) are deferred to Appendix B; reproducibility details (seeds, shared parameters) are in the code archive.
8.1 Claim 1: The Mini-Batch Covariance Is Identified by the Sampling Design
Before probing downstream consequences, we verify the covariance identification of Theorem 4.3 directly.
8.1.1 Experiment 1: Direct Covariance Identification (Probit DGP, Logistic Fit).
We generate n=50,000n=50{,}000 observations from a probit data-generating process and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$. Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}(\nabla_{\beta}\ell(\beta;X,Y)|_{\beta^{\star}})\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]|_{\beta^{\star}}$. The covariance-identification theorem nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx b^{-1}J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation. We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates). Figure 2 shows close agreement (relative Frobenius difference $\approx 0.05$): the identification holds even under misspecification, and mini-batch fluctuations are governed by the score covariance rather than by curvature alone.
Figure 2: Experiment 1: direct covariance identification under misspecification (probit DGP, logistic fit). Diagonal entries of the empirical mini-batch covariance vs. $b^{-1}\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.
8.2 Claim 2: The Lyapunov Geometry Is Visible in Low Dimensions
Once the identification is empirically in hand (Experiment 1), it is useful to visualize what the Lyapunov structure looks like in a two-dimensional setting before moving to the $d=10$ quantitative tests. In the fluctuation-scale OU surrogate of Theorem 5.7, the stationary covariance $\Sigma_{U}$ solves $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$, and Experiment 2 illustrates three consequences that are each most naturally displayed in $d=2$: the $1/b$ scaling of marginal variance, the redistribution of variance under curvature–noise misalignment, and the failure of trace-matched isotropic surrogates to reproduce directional structure. These are interpretive illustrations, not direct quantitative validation of SGD: the simulator integrates the time-changed OU process of (54) rather than running SGD iterates, and these subplots are included because a 2D visualization of the Lyapunov geometry aids intuition for the higher-dimensional quantitative tests that follow. The quantitative burden is carried by Experiments 3–6, which run direct SGD recursions.
8.2.1 Experiment 2: 2D Normalized OU Visualization of Lyapunov Structure.
For visualization we simulate a pedagogical normalized OU process chosen so that marginal variance scales as $1/b$ and is independent of $\eta$ at leading order: